A vessel is launched from Earth, carrying a popula
Answer #1

Since the can is travelling with the comparable speed of light(c) thus in this we need to use the concept of relativity. Then according to that

Time dilation or time taken by goats to die of old age will be t'= \frac{t}{\sqrt{1-\frac{v^2}{c^2}}}

a). Then if the speed is v= 0.845c then

Time taken, t'= \frac{15}{\sqrt{1-\frac{(0.845c)^2}{c^2}}}

  t'= \frac{15}{\sqrt{1-0.714}}= \frac{15}{0.534}= 28.08 yrs

Then distance travelled in this much time is

D'= vt'= 0.854 \times 3 \times 10^8 \times 28.08 \times 365 \times 24 \times 60 \times 60

D'= 2.268 \times 10^{17} m= (2.268 \times 10^{17} \times 1.057 \times 10^{-16}) ly

D'= 23.97 ly(ANS)

b). Now if the speed is v= 0.98c then

Time taken, t'= \frac{15}{\sqrt{1-\frac{(0.98c)^2}{c^2}}}

  t'= \frac{15}{\sqrt{1-0.96}}= \frac{15}{0.2}= 75 yrs

Then distance travelled in this much time is

D'= vt'= 0.98 \times 3 \times 10^8 \times 75 \times 365 \times 24 \times 60 \times 60

D'= 6.953 \times 10^{17} m= (6.953 \times 10^{17} \times 1.057 \times 10^{-16}) ly

D'= 73.49 ly(ANS)